OpenStudy (anonymous):

2x-4y=-14 -x-6y=-5 ,solve by substitution method. can someone show me step by step how to work it out please

OpenStudy (ajprincess):

2x-4y=-14-(1) -x-6y=-5-(2) Divide (2) by -1 Then u get x+6y=5 subtract 6y both sides x=5-6y-(3) Substitute this in (1) then u get 2(5-6y)-4y=-14 2*5-2*6y-4y=-14 10-12y-4y=-14 10+14=12y+4y 24=16y y=24/16 y=6/4 Substitute the value of y in (3) x=5-6*6/4 =(5*4-6*6)/4 =(20-36)/4 =-16/4 =-4 is it clear?

OpenStudy (anonymous):

yess but for the Y i saw that the answer is 3/2??

OpenStudy (ajprincess):

yup it is. y=6/4 =3/2

OpenStudy (anonymous):

i got that from the back of the book

OpenStudy (anonymous):

ok got it thank you! :)

OpenStudy (ajprincess):

yw.:)